OpenStudy (anonymous):

My answer is C What polynomial has roots of -4, 2, and 5 ? (3 points) Answer x3 - x2 - 22x + 40 x3 + x2 - 22x - 40 x3 + 3x2 - 18x - 40 x3 - 3x2 - 18x + 40

OpenStudy (anonymous):

( x +4) ( x - 2) ( x - 5) = 0

OpenStudy (anonymous):

can u multiply it

OpenStudy (anonymous):

4 x 2 x 5?

OpenStudy (anonymous):

= 40

OpenStudy (anonymous):

no multiply with the brackets

OpenStudy (anonymous):

so confused

OpenStudy (anonymous):

let me show u step by step

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(x + 4) (x (x - 5) - 2 ( x - 5) = ( x + 4) ( x^2 -5x -2x + 10)

OpenStudy (anonymous):

now x ( x^2 -5x -2x + 10) + 4 ( x^2 -5x -2x + 10)

OpenStudy (anonymous):

now multiply within the brackets

OpenStudy (anonymous):

x ( x^2 -5x -2x + 10) + 4 ( x^2 -5x -2x + 10) x^3 - 7x^2 + 10x + 4 x^2 - 28x + 40 x^3 - 3x^2 -18x + 40

OpenStudy (anonymous):

x^3 - 3x^2 -18x + 40 = 0....option D is the answer

OpenStudy (anonymous):

ohhh I'm starting to get the steps

OpenStudy (anonymous):

yes its very easy..